### v1

# : Adaptive Image Focus Optimization for Black-box MLLM

Abstract
Recent advancements in multimodal large language models (MLLMs) have broadened the scope of vision-language tasks, excelling in applications like image captioning and interactive question-answering. However, these models struggle with accurately processing visual data, particularly in tasks requiring precise object recognition and fine visual details. Stringent token limits often result in the omission of critical information, hampering performance. To address these limitations, we introduce , a novel visual prompting mechanism designed to enhance MLLM performance while preserving essential visual details within token limits. features three key innovations: a prompt-aware strategy that dynamically highlights relevant image regions, a spatial-preserving orchestration schema that maintains object integrity, and a budget-aware prompting method that balances global context with crucial visual details. Comprehensive evaluations across multiple datasets demonstrate that consistently outperforms baseline methods, achieving up to a improvement in accuracy while significantly reducing token consumption.
1 Introduction
Vision-language understanding has witnessed remarkable progress with multimodal large language models (MLLMs) such as GPT-4o, Gemini Pro, and Claude 3.5Β [13, 8]. While these black-box MLLMs have demonstrated impressive capabilities in controlled settingsβeven being touted as having PhD-level skillsβour large-scale investigation reveals a critical hallucination problem: their inherent small object blindness and limited visibility lead to catastrophic failures in real-world visual reasoning tasks.
Through systematic analysis, we identify two fundamental limitations that create this problem. First, despite their sophisticated architectures, current black-box MLLMs process images with uniform attention, lacking the human-like ability to focus on relevant regions while maintaining spatial awareness. As shown in FigureΒ 1, this limitation leads to surprising failures in entry-level tasks like object countingβa problem that persists across different models and prompting strategies. Our investigation reveals that this is not merely an implementation issue but a fundamental limitation of how these models process visual information.
Second, we identify a critical tension between token budgets and visual fidelity in black-box MLLMs. While commodity black-box MLLM has treated token limitations, typically fixed token counts, as a simple engineering constraint to ensure computational efficiency and user fairnessΒ [4], our analysis reveals it as a fundamental resource allocation problem. As demonstrated in FigureΒ 2, the standard approach of uniform downsampling leads to catastrophic loss of fine detailsβa problem that cannot be solved by better downsampling algorithms or prompt engineering alone.
This investigation leads us to formulate a novel problem in visual prompting: How can we achieve region-aware visual processing within the strict token constraints of black-box MLLMs? This question challenges the conventional view that visual prompting is merely a simple input formatting task. Instead, we show that effective visual prompting requires solving three interconnected challenges: Region Selection: How to identify and prioritize task-relevant regions without access to model internals; Spatial Preservation: How to maintain structural relationships while maximizing detail preservation; Budget Optimization: How to allocate limited tokens between global context and local details.
Based on these insights, we propose , a visual prompting framework that addresses these challenges through three technical innovations:
β’
A prompt-aware visual emphasis mechanism that enables selective attention without modifying model architecture
β’
A spatial-preserving orchestration schema that maintains global context while preserving local details
β’
A budget-aware region selection strategy that optimally allocates tokens across image regions
In a comprehensive evaluation across datasets such as VstarΒ [28], CVBenchΒ [24], and RealworldQAΒ [29], consistently outperformed baseline methods. Notably, in the Vstar dataset, -Patches achieved a 26.9% accuracy improvement over the baseline, while in RealWorldQA, -Adaptive outperformed the baseline by 12.1%. Alongside these accuracy gains, significantly reduced token usage. For example, on the TerraIncognita dataset, achieved 6.4% higher accuracy with a 67% reduction in token consumption compared to the baseline. These results confirm that not only addresses the limitations of visual processing in black-box MLLMs but also enhances efficiency. Moreover, across APIs like GPT-4o111https://platform.openai.com/, Gemini-1.5Pro222https://gemini.google.com/, and Claude-3.5-Sonnet333https://anthropic.com/, demonstrated consistent improvements in both accuracy and token efficiency, solidifying its potential to optimize MLLM performance in real-world, high-resolution applications.
The contribution of this work can be summarized as follows: (1) We conduct a detailed investigation of GPT-4oβs image prompting strategy, exposing key limitations in its handling of visual inputs. (2) We introduce , a novel mechanism that addresses the challenge of preserving visual detail in black-box MLLMs while adhering to token constraints. (3) We present extensive experimental results across multiple datasets, demonstrating that achieves significant improvements in both accuracy and token efficiency, offering valuable insights into enhancing multimodal processing in constrained environments.
2 Pilot Experiments
One of the primary challenges faced by black-box MLLMs is their inability to process visual inputs efficiently, leading to diminished accuracy in fine-grained visual tasks. Models like GPT-4o often struggle with recognizing detailed or occluded objects, particularly when dealing with complex images. These limitations are compounded by token constraints, which restrict the amount of image data that can be processed in a single prompt. Furthermore, according to the Vision pricing calculator444https://openai.com/api/pricing/, GPT handle images by resizing and splitting them into basic units of 512Γ512 pixels. Each of these units corresponds to 170 tokens. This method of processing not only imposes a strict limit on the image resolution but also increases the computational overhead due to the additional tokens generated from splitting. As a result, vital visual details may be lost when images are downsampled or resized to fit within the token limits, leading to poor performance on tasks that require precise visual grounding.
| Method | Accuracy |
| Unaltered Input | 57% | 955 | ||
| Image Crop | 58% | 270 | ||
| Zoomed Crop | 64% | 270 |
To evaluate this issue, we conducted a series of pilot experiments using GPT-4o-0513 on the Vstar-Bench dataset. This dataset challenges MLLMs to accurately identify detailed objects within high-resolution images, making it an ideal test for the modelβs capacity to handle fine-grained visual information. The experiments compared three different image processing strategies: (1) an unprocessed prompt (Unaltered Input), where the image is fed to the model in its original form; (2) a prompt where the image is cropped to focus on the target object (Image Crop); and (3) a prompt where the cropped image is further enlarged to emphasize the most relevant visual features (Zoomed Crop). Both the Image Crop and Zoomed Crop methods were constrained to fit within GPT-4oβs patch size limit of 512x512 pixels.
As shown in TableΒ 1, the Zoomed Crop method significantly outperformed the others, achieving an accuracy of 0.76 with a token usage of 270. In comparison, the Unaltered Input method, despite processing the entire image, only achieved an accuracy of 0.64 while consuming 955 tokens. Similarly, the Image Crop method, although reducing the token count to 270, did not yield any improvement in accuracy compared to the unprocessed input.
These results highlight a fundamental problem in current black-box MLLMs: they fail to efficiently manage the trade-off between image resolution and token constraints. In cases like those presented by the Vstar-Bench dataset, where fine-grained visual information is critical, processing unaltered high-resolution images leads to excessive token consumption without improving accuracy. While the Image Crop method reduces token usage, it fails to improve performance because simply cropping an image without emphasizing key details does not provide sufficient context for the model to interpret the visual input accurately.
The superior performance of the Zoomed Crop method underscores the importance of vision enhancement techniques in black-box MLLMs. By focusing on the most relevant portions of the image, Zoomed Crop preserves critical details while remaining within token limits, enabling the model to interpret detailed visual inputs more effectively. This approach resolves a common issue faced by black-box MLLMs, where downscaling or cropping images to meet token requirements often leads to a loss of essential information, reducing the overall effectiveness of the model.
Our experiments reveal that without adaptive techniques like Zoomed Crop, black-box MLLMs struggle to process high-resolution images efficiently, limiting their performance on tasks that require precise visual recognition. These findings demonstrate the necessity of vision enhancement strategies to address the inherent limitations of token-constrained MLLMs.
3 Related Work
3.1 Multimodal LLMs: Open-Source and Black-Box Models
The integration of visual and textual modalities in large language models (LLMs) has led to significant advancements in multimodal models (MLLMs) like GPT-4o, Gemini Pro and Claude3-Sonnet. These models rely on effective visual encoding strategies to bridge the gap between language and vision. Approaches such as CLIPΒ [31] align visual and language embeddings through contrastive learning, while models like FlamingoΒ [1] and BLIP-2Β [6] use cross-attention mechanisms or pretraining modules to link vision encoders with LLMs. However, these methods often rely on fixed low-resolution inputs (e.g., 224x224), limiting their ability to process high-resolution images or non-standard aspect ratiosΒ [16], which hampers performance on fine-grained tasks such as OCR and small object detection
In contrast, open-source multimodal modelsΒ [15, 30, 36, 12, 37] allow for architectural modifications and fine-tuning to accommodate any-resolution inputs. However, black-box MLLMs such as GPT-4o and Gemini Pro, which impose strict token limits for computational efficiency, require alternative solutions. The need to downsample or crop images to meet these constraints often results in the loss of crucial visual details, particularly in tasks requiring detailed visual understanding. While position embedding interpolationΒ [2, 26, 19, 9, 5] and patch-based croppingΒ [30, 12] widely adpoted in open-soure models offer promising directions for any aspect ratio and any-resolution image processing, they are not applicable to black-box models, where architectural changes and extra training/fine-tuning are not permitted.
3.2 Object Detection
Traditional object detection models, such as Faster R-CNNΒ [22] and YOLOΒ [21], effectively identify and localize objects within predefined categories. However, they struggle with open-set scenarios, where novel objects not seen during training need to be detected.
Recent advances address this limitation through open-set detection models that leverage natural language processing. For instance, OV-DETRΒ [34] integrates CLIP with object detection to generate category-specific bounding boxes from textual prompts, enabling detection in open-world settings. Similarly, GLIPΒ [14] reframes detection as a grounding problem, improving alignment between visual regions and textual descriptions. DetCLIPΒ [32] extends this further using pseudo labels from large-scale captioning datasets, enhancing generalization. Grounding DINOΒ [17], built on the DETR frameworkΒ [3], also advances open-set detection through natural language integration.
In addition, SAMΒ [11] and SAM-2Β [20] offer zero-prompt or minimal-prompt segmentation for arbitrary objects but lack robust text-prompt handling. EVF-SAMΒ [35] overcomes this by extending SAMβs capabilities to better manage complex text-based object segmentation.
By incorporating these models, enhances its ability to dynamically detect and emphasize regions of interest (RoIs), enabling black-box MLLMs to focus on the most relevant visual content without losing critical details, which is essential for maintaining high performance across varied resolutions.
4 Method Overview
Inspired by the observation derived from our pilot experiments, we propose , a comprehensive visual prompting mechanism designed to effectively address the loss of detail in images that occurs during the naive resizing process in current black box multimodal LLMs, such as GPT-4 and Gemini 1.5. As illustrated in FigureΒ 3, our mechanism comprises three key components: (1) A prompt-aware visual emphasizer that allocates high-fidelity image slices based on prompt texts to facilitate efficient and focused visual encoding; (2) A spatial-preserving encoding schema that consolidates the collected image slices while maintaining their relative spatial positions to create a condensed visual input; (3) A budget-aware prompting strategy that maximizes the accuracy of results obtained from the black box models while fits the budget requirement from users.
4.1 Prompt-aware Visual Emphasizer
The prompt-aware visual emphasizer utilizes a multi-scale emphasizing strategy to prioritize image slices that are most relevant to the input prompts. By analyzing the semantic content of the prompts, this component dynamically selects and enhances specific regions of the image at varying resolutions. This approach not only enriches the contextual information available to the model but also mitigates the adverse effects of losing critical details during the resizing process.
Prompt Tokenization Prompt tokenization is a critical first step in which input prompts are parsed into meaningful tokens. This process segments the prompt into components that can be easily analyzed for semantic relevance. Specifically, the prompt is divided into structural components, and our focus is on processing the relevant sections that contribute directly to visual emphasis.
To enhance the extraction of semantically relevant tokens, we apply advanced natural language processing (NLP) techniques. First, we use the NLTK library555https://www.nltk.org/ to remove stopwords, reducing noise and ensuring that the modelβs attention remains on the most critical visual elements. By eliminating these non-essential words, we concentrate on key terms that directly influence the visual emphasis.
In addition to basic stopword removal, we utilize dependency parsingΒ [23, 7] to analyze the syntactic structure of the prompt. This deeper analysis identifies core entities and relationships, such as subject-object pairs and action verbs, which are crucial for interpreting the userβs intent. By focusing on these core semantic elements, we ensure that the visual emphasis aligns precisely with the underlying meaning of the prompt.
Finally, we strip away any irrelevant formatting or non-content-related details, allowing the visual emphatizer to focus solely on the essential information. This multi-layered tokenization approach ensures an optimal match between the tokenized prompt and the image features selected for emphasis.
Multi-Scale Emphasizing Algorithm: Given a key object term extracted from the text prompt, the Multi-Scale Emphasizing AlgorithmΒ 1 utilizes a state-of-the-art object detection model to localize the corresponding object in the image prompt. In our experiments, we primarily employ GroundingDINOΒ [17] as our localization model.
The encoder in such models typically downsamples the input image to a resolution of or , potentially resulting in information loss when localizing the target object at a coarse granularity. To address this limitation, we propose a Multi-Scale Emphasizing Algorithm that processes the original image at multiple resolutions. The algorithm divides the input image into patches at various granularities, e.g. , , and beyond. For each generated patch, we apply the object detection model to localize the target object. The algorithm retains bounding boxes returned by the model that exceed a predefined confidence threshold. These high-confidence bounding boxes collectively form the output of our algorithm, providing a comprehensive multi-scale representation of the target objectβs location.
4.2 Spatial-preserving Orchestration Schema
Building upon the Multi-Scale Emphasizing Algorithm, we introduce a Spatial-preserving Orchestration Schema to maintain the structural integrity of the image during the encoding process. This schema filters the bounding boxes obtained from the Multi-Scale Emphasizing Algorithm and ensures that the relative positions of the selected image slices are preserved, facilitating a more faithful representation of the original image layout and enabling coherent reconstruction when processed by the multimodal LLM. To refine the selection of bounding boxes, we implement a Non-Maximum Suppression (NMS) based slice filtering method. NMS is employed to eliminate redundant and overlapping slices, retaining only the most salient features that align with the prompt. The process works as described in AlgorithmΒ 2.
By setting an appropriate threshold for the Intersection of Union (IoU) of bounding boxes around the selected regions, we ensure that only the highest-quality slices are retained for the encoding process. This filtering step enhances computational efficiency by reducing the number of slices to be processed and improves the clarity and relevance of the visual information provided to subsequent stages of the model.
The resulting set of filtered slices are then orchestrated to preserve their original relative positions within the image. This orchestration process involves the following steps: Slice Extraction: For each bounding box in the filtered set , we extract the corresponding image slice from the original image. Blank Image Creation: We create a new blank image with the same dimensions as the original image. Slice Placement: We place each extracted slice onto the blank image at its original position, leaving the rest of the image blank. Image Shrinking: The resulting image, containing only the selected slices in their original positions with the rest left blank, is then shrunk to a predetermined size while maintaining its aspect ratio.
4.3 Budget-aware Prompting Strategy:
Our approach incorporates a sophisticated budget-aware prompting strategy that optimizes the allocation of token budget for image processing. This strategy begins with a user-specified total token budget , allowing for customization based on specific task requirements or computational constraints. We propose four varieties of to accommodate different budget scenarios and task requirements:
β’
-Local(β£): This variant utilizes only the spatial-preserving schema to consolidate all focused image slices into a single image patch(β£ in FigureΒ 3). It is optimal for scenarios with very limited token budgets, prioritizing the most relevant visual information.
β’
-Adaptive (β£ + β’ ): This approach dynamically includes a global view of the original image if the cropped portion falls below a certain threshold . This allows the MLLM to better understand the overall scene context when the budget permits, while still focusing on key areas of interest.
β’
-Global (β£ + β’): This variant assigns a global view to all images, regardless of the specific regions of interest. It is suitable for tasks that require consistent overall context and when the token budget is sufficient to include both global and local information.
β’
-Patches(β‘ + β’): This is the most token-intensive approach, assigning each image slice its own patch without spatial preservation, along with a global view. It provides the most detailed information but requires the largest token budget.
The selection among these varieties depends on the user-specified budget and the nature of the task. For each variant, the number of high-resolution slices or patches is calculated based on the available budget and the token cost per slice or patch. These slices are selected from the output of our Multi-Scale Emphasizing Algorithm, prioritizing based on their relevance to the key term of text prompts. To present the methods more clearly and vividly, we refer to FigureΒ 4, which outlines the methodology, and FigureΒ 5, which showcases a specific case study.
5 Experiments
| Vstar | CVBench-2D | CVBench-3D | RealworldQA | SQA-I | MMVP | MMMU | HR-4K | HR-8K | |
| Raw | 56.5%/955 | 68.5%/428 | 78.2%/895 | 67.6%/998 | 87.3%/353 | 83.3% /270 | 68.4%/608 | 50.6%/1105 | 46.8%/1105 |
| Resize | 41.9%/270 | 66.3%/270 | 75.2%/270 | 61.1%/270 | 86.8%/270 | 83.3%/270 | 62.9%/270 | 35.8%/270 | 33.4%/270 |
| -Local | 67.1%/270 | 72.4%/270 | 86.2%/270 | 72.4%/270 | 88.3%/270 | 87.1%/270 | 59.8%/270 | 58.8%/270 | 57.7%/270 |
| -Adaptive | 67.5%/419 | 72.9%/374 | 87.9%/408 | 74.7%/362 | 91.1%/308 | 88.7%/351 | 61.6%/312 | 60.8%/331 | 59.3%/324 |
| -Global | 67.6%/540 | 73.1%/540 | 88.3%/540 | 75.3%/540 | 92.3%/540 | 88.9%/540 | 67.3%/540 | 61.3%/540 | 59.8%/540 |
| -Patches | 71.7%/1029 | 74.6%/709 | 85.8%/1113 | 75.8%/997 | 92.8%/727 | 88.4%/726 | 68.9%/841 | 60.4%/713 | 58.9%/875 |
| Method | Accuracy | Tokens | Latency | Money Cost($10-e3) | ||||
| Zero-Shot | 15-Shot | Zero-Shot | 15-Shot | Zero-Shot | 15-Shot | Zero-Shot | 15-Shot | |
| Raw | 78% | 84% | 963 | 13488 | 4.8s | 18.7s | 4.815 | 67.44 |
| Resize | 61% | 74% | 255 | 4080 | 2.9s | 7.5s | 1.275 | 20.4 |
| Low-Detail | 60% | 70% | 85 | 1360 | 2.1s | 6.5s | 0.425 | 6.8 |
| -Adaptive | 83% | 88% | 315 | 5112 | 3.1s | 9.8s | 1.575 | 25.56 |
| API | Method | Vstar | CVBench-2D | RealworldQA | MMVP |
| GPT-4o | Raw | 56.5% | 68.5% | 67.6% | 83.3% |
| 71.7% | 74.6% | 75.8% | 88.9% | ||
| Gemini- 1.5Pro | Raw | 53.1% | 65.4% | 64.0% | 79.8% |
| 70.4% | 73.2% | 73.9% | 87.8% | ||
| Claude- 3.5-Sonnet | Raw | 51.8% | 66.7% | 61.0% | 80.2% |
| 69.7% | 72.8% | 74.1% | 87.2% |
| Method | Model | Prompt Strategy | Vstar | CVBench-2D | CVBench-3D | RealworldQA | SQA-I | MMVP | MMMU |
| Default | EVF-SAM | Local | 57.1% | 67.6% | 79.9% | 68.5% | 87.8% | 84.0% | 55.3% |
| Adaptive | 57.5% | 71.3% | 82.5% | 72.1% | 88.3% | 85.3% | 55.9% | ||
| Global | 57.8% | 72.1% | 83.0% | 72.4% | 88.8% | 87.3% | 56.8% | ||
| Patches | 57.1% | 72.7% | 83.8% | 72.1% | 86.8% | 84.7% | 56.1% | ||
| Ground Dino | Local | 58.1% | 70.6% | 82.5% | 71.5% | 85.3% | 84.3% | 55.6% | |
| Adaptive | 58.3% | 71.5% | 83.9% | 73.1% | 90.1% | 85.6% | 56.3% | ||
| Global | 58.3% | 71.8% | 84.8% | 73.4% | 90.3% | 87.8% | 57.0% | ||
| Patches | 58.8% | 70.6% | 83.1% | 72.6% | 90.9% | 87.7% | 56.6% | ||
| Multi- Resolution | EVF-SAM | Local | 58.4% | 69.2% | 84.0% | 72.5% | 88.3% | 84.7% | 57.1% |
| Adaptive | 58.5% | 71.3% | 85.3% | 73.1% | 91.1% | 86.8% | 58.7% | ||
| Global | 58.4% | 72.1% | 85.8% | 73.4% | 91.8% | 87.8% | 59.6% | ||
| Patches | 60.2% | 71.3% | 85.7% | 73.1% | 91.3% | 86.8% | 59.2% | ||
| Ground Dino | Local | 63.6% | 71.8% | 82.6% | 70.2% | 88.8% | 85.1% | 56.8% | |
| Adaptive | 64.4% | 71.8% | 84.3% | 70.6% | 90.3% | 86.5% | 58.5% | ||
| Global | 66.4% | 72.2% | 84.7% | 71.4% | 91.8% | 86.9% | 59.3% | ||
| Patches | 66.2% | 72.6% | 83.6% | 70.2% | 92.1% | 86.9% | 58.8% | ||
| Multi-Scale | EVF-SAM | Local | 63.7% | 72.1% | 85.2% | 70.4% | 85.3% | 85.7% | 57.5% |
| Adaptive | 64.3% | 72.4% | 86.1% | 70.6% | 90.1% | 87.6% | 58.5% | ||
| Global | 64.3% | 72.8% | 87.9% | 71.7% | 90.3% | 88.8% | 59.9% | ||
| Patches | 67.2% | 73.7% | 87.1% | 73.0% | 90.9% | 88.0% | 59.7% | ||
| Ground Dino | Local | 67.1% | 72.4% | 86.2% | 72.4% | 88.3% | 87.1% | 57.7% | |
| Adaptive | 67.5% | 72.9% | 87.9% | 74.7% | 91.1% | 88.7% | 59.3% | ||
| Global | 67.6% | 73.1% | 88.3% | 75.3% | 92.3% | 88.9% | 59.8% | ||
| Patches | 71.7% | 74.6% | 85.8% | 75.8% | 92.8% | 88.4% | 58.9% |
In this section, we evaluate the performance of through a series of experiments designed to test its ability to improve token efficiency and preserve visual fidelity across different black-box MLLM. Specifically, we aim to answer the following questions: (i) Accuuracy: Does improve accuracy across different black-box MLLMs on image-related tasks? (ii) Efficiency: How does perform compared to baseline methods in terms of both accuracy and token efficiency? (iii) Component Contribution: What is the impact of key components in , such as multi-scale vision emphasize and the budget-aware prompt strategy?
5.1 Setup
Assessment and Datasets We evaluated our system on a series of challenging multimodal tasks, using commercial black-box MLLMs for applications ranging from visual-language reasoning to image understanding and question answering. The experiments were conducted on a variety of different public datasets, including:
1) VstarΒ [28]: A benchmark dataset focused on image classification, used to evaluate fine-grained visual recognition capabilities in object detection and classification tasks.
2) CVBenchΒ [24]: Contains 2 sub-category, and , respectively, representing two-dimensional and three-dimensional visual image, respectively, to evaluate the performance of the model when processing images of different dimensions, especially the understanding ability in complex scenes.
3) RealworldQAΒ [29]: Used to test the multimodal question answering performance of the model in real-world scenarios, involving cross-language and cross-image information processing.
4) MMVPΒ [25]: A validation set for multimodal visual processing, designed to evaluate the comprehensive understanding of models for complex visual scenes.
5) ScienceQAΒ [18]: A multimodal scientific question-answering dataset featuring multiple-choice questions across a diverse range of science topics.
6) MMMUΒ [33]: The validation part of a new benchmark, which designed to evaluate the performance of multimodal models on multidisciplinary tasks that require university-level subject knowledge and deliberate reasoning.
7) HRΒ [27]: A high-resolution multimodal benchmark consisting of 4K and 8K images and corresponding questions.
Models We employed three black-box MLLMsβGPT-4o-0513, Claude-v3-Sonnet, and Gemini-Proβaccessed via their respective APIs (OpenAI, Claude, Google). Across all experiments, we set the temperature to 0 and used greedy decoding for consistency, optimizing the stability of outputs. NMS was applied with a confidence score threshold of 0.8 to filter irrelevant regions from high-resolution images.
Metrics We used classification accuracy across all examples as the primary evaluation metric. Additionally, we compared token usage for each model configuration to evaluate the efficiency improvements offered by .
Baselines We compare against the following baseline methods:
1) Raw: This baseline feeds MLLM the unmodified prompt, with no adjustments made to the image.
2) Resize: Here, images larger than 512x512 pixels are resized to fit within the GPT-4oβs patch limit, while smaller images remain unchanged.
5.2 Main results
TableΒ 2 compares the performance of against baseline methods across various datasets using GPT-4o. The results show that , particularly in its Patches and Adaptive versions, consistently outperformed baseline approaches in terms of accuracy while maintaining lower token usage than the Raw method. consistently outperformed baseline methods, showing accuracy improvements up to 26% across multiple tasks. For example, in the Vstar dataset, -Patches achieved an accuracy of 0.717, compared to 0.565 using the Raw baseline, marking a 26.9% improvement. In RealworldQA, which demands complex multimodal reasoning, -Adaptive achieved 0.758 accuracy, outperforming the 0.676 accuracy of the Raw method by 12.1%. These results highlight that effectively preserves fine-grained visual details, enabling improved object recognition and image understanding across real-world tasks, where precise detail retention is crucial.
We further evaluated on Claude-3.5-Sonnet and Gemini-1.5Pro to assess its generalizability across different black-box MLLMs. TableΒ 4 shows that demonstrated robust performance across different black-box MLLMs, for example, achieved an accuracy of 0.704 on Vstar, compared to 0.531 with the Raw baseline, marking a 32.6% improvement. Similarly, in Claude-3.5-Sonnet, outperformed the baseline by 34.5% on RealworldQA, improving from 0.610 to 0.741. These results suggest that can consistently enhance performance across different architectures, making it a versatile tool for various MLLM-based applications.
A key contribution of is its ability to reduce token consumption while maintaining or improving accuracy. As shown in TableΒ 3, consistently delivers both token efficiency and performance improvements across various benchmark datasets. For instance, on the TerraIncognita dataset, guided by ManyICLΒ [10], achieves 0.83 accuracy using 315 tokens, compared to the Raw baselineβs 0.78 accuracy with 963 tokensβa 67% reduction in token usage while improving performance by 6.4%. Additionally, reduces latency, making it practical for real-time applications. In the TerraIncognita zero-shot setting, lowered latency from 4.8s to 3.1s, a 35.4% reduction without sacrificing accuracy. This makes highly suitable for tasks like autonomous driving, and real-time visual analytics, where both token efficiency and reduced latency are critical.
5.3 Findings
Here we analyze why the -Patches version underperforms compared to -Global and even the Local version on certain datasets. For example, on the dataset, the accuracy of the Patches version is 0.025 lower than the Global version and 0.04 lower than the Local version. Similarly, on the MMVP dataset, the Patches version falls short by 0.005 compared to the Global version and by 0.003 compared to the Adaptive version. Given that these results are averaged over multiple measurements, and accounting for model fluctuations, we hypothesize that this performance drop occurs because the Patches version treats each RoI as an independent image and provides them separately to the MLLM. When there are too many RoIs, the model may fail to capture or integrate some of them, leading to a drop in accuracy.
5.4 Ablation study
To further understand the impact of the individual components within , we conducted an ablation study focusing on two key variants: 1) with multi-scale emphasize: Compared with the commonly used multi-resolution and directly use, multi-scale visual emphasis is used to identify and emphasize RoI in the image. 2) with different models.
To further investigate the contributions of βs components, we conducted an ablation study (TableΒ 5). The results demonstrate that the combination of multi-scale visual emphasis and Patches prompt strategies delivers the best performance across all most datasets. Comparing different vision emphasis models, such as EVF-SAM and Ground Dino, further highlights the effectiveness of . Despite differences in model capabilities, both show accuracy improvements across datasets. Additionally, when comparing different emphasis methodsβDefault, Multi-Resolution, and Multi-Scaleβthe Multi-Scale method consistently outperformed Multi-Resolution. We hypothesize that, while Multi-Scale crops images and may split objects, its pyramid-shaped multi-recall strategy compensates for this by enhancing recall. In contrast, although Multi-Resolution maintains object integrity, adjusting resolution disrupts the modelβs performance, likely because most models are trained on fixed-size inputs, and changing the resolution weakens their inherent capabilities.
6 Conclusion
In this paper, we introduced , a novel visual prompting mechanism designed to overcome the limitations of black-box MLLMs in processing images while adhering to token constraints. Our approach effectively balances the need to capture essential visual details without exceeding token budgets, a challenge commonly encountered in existing models like GPT-4o and Gemini Pro.
Through a comprehensive evaluation across datasets such as Vstar and RealWorldQA, demonstrated significant improvements, particularly in fine-grained visual tasks. Our results show that -Patches achieved a 26.9% accuracy gain over baseline methods in Vstar, and -Adaptive provided a 12.1% improvement in RealWorldQA. These gains were achieved while drastically reducing token usage, with delivering 6.4% higher accuracy in the TerraIncognita using 67% fewer tokens.
Although this work primarily focuses on improving the efficiency of visual processing in black-box MLLMs, another potential issue that arises in real-world applications is communication cost. For example, transferring large images from edge devices (e.g., wearable cameras or glasses) to cloud servers can be expensive in terms of bandwidth, latency, and energy consumption. While is designed to reduce token usage, its application in minimizing data transmission costs is an area that could be explored in future work.
As part of future research, we plan to investigate how could be adapted for edge ML applications, enabling local processing on devices such as wearable cameras. This would allow for more efficient handling of visual inputs at the edge, reducing the need for extensive data transfers to the cloud. We aim to measure latency, power consumption, and the overall impact on system performance to assess the feasibility of applying in these scenarios.
In summary, offers a practical solution for enhancing visual processing in constrained MLLMs and opens up new directions for future exploration.

### v2

@@
-: Adaptive Image Focus Optimization for Black-box MLLM
+: Adaptive Visual Token Allocation for Black-box MLLM
 Abstract
-Recent advancements in multimodal large language models (MLLMs) have broadened the scope of vision-language tasks, excelling in applications like image captioning and interactive question-answering.
-However, these models struggle with accurately processing visual data, particularly in tasks requiring precise object recognition and fine visual details.
-Stringent token limits often result in the omission of critical information, hampering performance.
-To address these limitations, we introduce , a novel visual prompting mechanism designed to enhance MLLM performance while preserving essential visual details within token limits.
-features three key innovations: a prompt-aware strategy that dynamically highlights relevant image regions, a spatial-preserving orchestration schema that maintains object integrity, and a budget-aware prompting method that balances global context with crucial visual details.
-Comprehensive evaluations across multiple datasets demonstrate that consistently outperforms baseline methods, achieving up to a improvement in accuracy while significantly reducing token consumption.
+Multimodal large language models (MLLMs) such as GPT-4o, Gemini Pro, and Claude 3.5 have enabled unified reasoning over text and visual inputs, yet they often hallucinate in real-world scenariosβespecially when small objects or fine spatial context are involved.
+We pinpoint two core causes of this failure: the absence of region-adaptive attention and inflexible token budgets that force uniform downsampling, leading to critical information loss.
+To overcome these limitations, we introduce , a visual prompting framework that delivers token-efficient, detail-preserving image representations for black-box MLLMs.
+integrates (1) a prompt-aware emphasis module to highlight semantically relevant regions, (2) a spatial-preserving orchestration schema to maintain object relationships, and (3) a budget-aware strategy to adaptively allocate tokens between global context and local details.
+Extensive experiments on nine benchmarks and three commercial MLLMs demonstrate that boosts accuracy by up to 27% while cutting image token usage by up to 67%.
+Our approach establishes a principled methodology for robust, resource-aware multimodal understanding in settings where model internals are inaccessible.
 1 Introduction
-Vision-language understanding has witnessed remarkable progress with multimodal large language models (MLLMs) such as GPT-4o, Gemini Pro, and Claude 3.5Β [13, 8].
-While these black-box MLLMs have demonstrated impressive capabilities in controlled settingsβeven being touted as having PhD-level skillsβour large-scale investigation reveals a critical hallucination problem: their inherent small object blindness and limited visibility lead to catastrophic failures in real-world visual reasoning tasks.
-Through systematic analysis, we identify two fundamental limitations that create this problem.
-First, despite their sophisticated architectures, current black-box MLLMs process images with uniform attention, lacking the human-like ability to focus on relevant regions while maintaining spatial awareness.
-As shown in FigureΒ 1, this limitation leads to surprising failures in entry-level tasks like object countingβa problem that persists across different models and prompting strategies.
-Our investigation reveals that this is not merely an implementation issue but a fundamental limitation of how these models process visual information.
-Second, we identify a critical tension between token budgets and visual fidelity in black-box MLLMs.
-While commodity black-box MLLM has treated token limitations, typically fixed token counts, as a simple engineering constraint to ensure computational efficiency and user fairnessΒ [4], our analysis reveals it as a fundamental resource allocation problem.
-As demonstrated in FigureΒ 2, the standard approach of uniform downsampling leads to catastrophic loss of fine detailsβa problem that cannot be solved by better downsampling algorithms or prompt engineering alone.
-This investigation leads us to formulate a novel problem in visual prompting: How can we achieve region-aware visual processing within the strict token constraints of black-box MLLMs?
-This question challenges the conventional view that visual prompting is merely a simple input formatting task.
-Instead, we show that effective visual prompting requires solving three interconnected challenges: Region Selection: How to identify and prioritize task-relevant regions without access to model internals; Spatial Preservation: How to maintain structural relationships while maximizing detail preservation; Budget Optimization: How to allocate limited tokens between global context and local details.
-Based on these insights, we propose , a visual prompting framework that addresses these challenges through three technical innovations:
+Vision-language understanding has emerged as a central challenge in multimodal artificial intelligence, with broad applications ranging from robotics and autonomous driving to scientific diagram analysis and human-computer interaction.
+Recent advances in multimodal large language models (MLLMs), such as GPT-4o, Gemini Pro, and Claude 3.5, have significantly pushed the boundaries of this field by enabling unified reasoning over text and visual inputsΒ li2024evcap; gu2024anomalygpt.
+These models have demonstrated impressive performance on a range of benchmarks and are increasingly perceived as possessing near-human or even PhD-level capabilities in controlled environments.
+However, despite these promising results, our large-scale empirical study uncovers a critical limitation that remains underexplored: MLLMs exhibit a pronounced hallucination tendency in real-world visual reasoning tasks, largely due to their inherent difficulty in perceiving small objects and limited field-of-view awareness.
+This βsmall object blindnessβ can lead to severe and often undetected reasoning failures, raising concerns about the deployment of MLLMs in real-world visual reasoning scenarios.
+To investigate the root causes of hallucination in MLLMs, we conduct a systematic analysis and uncover two fundamental limitations in the visual processing pipeline of current black-box MLLMs.
+First, these models lack region-adaptive attention mechanisms, resulting in uniform spatial processing that disregards task-relevant visual saliency.
+Despite their architectural complexity, black-box MLLMs fail to emulate the human visual systemβs ability to focus selectively while preserving contextual structure.
+As illustrated in FigureΒ 1, this limitation leads to persistent failure even in elementary tasks such as object counting.
+Notably, the issue is consistent across different model families and prompting strategies, indicating a fundamental limitation rooted in the design of their visual tokenization and attention mechanisms, rather than implementation-specific flaws.
+Second, we identify a critical tension between token budget constraints and visual fidelity.
+Black-box MLLMs are constrained by fixed token limits for both textual and visual inputsβan engineering choice originally motivated by inference efficiency and fairness considerationsΒ (chen2024spatialvlm).
+However, our analysis reveals that token budgeting in vision tasks is a core computational bottleneck that directly impacts perception granularity.
+As shown in FigureΒ 2, the commonly adopted strategy of uniform image downsampling leads to substantial loss of local detail, especially for small or spatially distributed objects.
+Crucially, we demonstrate that neither prompt engineering nor improved resizing algorithms can compensate for this representational loss.
+These observations motivate us to formally define a new problem in visual prompting for MLLMs: How can region-aware, spatially coherent visual inputs be constructed under strict token constraints in black-box inference settings?
+We argue that solving this problem requires a paradigm shift from heuristic prompt design to principled visual token orchestration.
+We decompose this problem into three interdependent sub-tasks: (1) Region Selection β identifying and prioritizing task-relevant regions without model internals; (2) Spatial Preservation β retaining spatial relations and object continuity across selected regions; and (3) Budget-Aware Prompting β balancing the allocation of tokens between global context and fine-grained details.
+To address these challenges, we present , a visual prompting framework designed for token-efficient, region-aware visual representation. comprises three integrated components:
 β’
-A prompt-aware visual emphasis mechanism that enables selective attention without modifying model architecture
+A prompt-aware visual emphasis module, which selectively highlights salient regions by manipulating image input structure, enabling focused attention without altering the model;
 β’
-A spatial-preserving orchestration schema that maintains global context while preserving local details
+A spatial-preserving orchestration scheme, which ensures that contextual coherence and relative object positioning are preserved across split or patched inputs;
 β’
-A budget-aware region selection strategy that optimally allocates tokens across image regions
-In a comprehensive evaluation across datasets such as VstarΒ [28], CVBenchΒ [24], and RealworldQAΒ [29], consistently outperformed baseline methods.
-Notably, in the Vstar dataset, -Patches achieved a 26.9% accuracy improvement over the baseline, while in RealWorldQA, -Adaptive outperformed the baseline by 12.1%.
-Alongside these accuracy gains, significantly reduced token usage.
-For example, on the TerraIncognita dataset, achieved 6.4% higher accuracy with a 67% reduction in token consumption compared to the baseline.
-These results confirm that not only addresses the limitations of visual processing in black-box MLLMs but also enhances efficiency.
-Moreover, across APIs like GPT-4o111https://platform.openai.com/, Gemini-1.5Pro222https://gemini.google.com/, and Claude-3.5-Sonnet333https://anthropic.com/, demonstrated consistent improvements in both accuracy and token efficiency, solidifying its potential to optimize MLLM performance in real-world, high-resolution applications.
-The contribution of this work can be summarized as follows: (1) We conduct a detailed investigation of GPT-4oβs image prompting strategy, exposing key limitations in its handling of visual inputs.
-(2) We introduce , a novel mechanism that addresses the challenge of preserving visual detail in black-box MLLMs while adhering to token constraints.
-(3) We present extensive experimental results across multiple datasets, demonstrating that achieves significant improvements in both accuracy and token efficiency, offering valuable insights into enhancing multimodal processing in constrained environments.
+A budget-aware region selection strategy, which dynamically allocates token capacity based on content density and task relevance.
+We evaluate on a suite of diverse benchmarks, including Vstar(vstar), CVBench(tong2024cambrian1), and RealWorldQAΒ (xAI2024), which collectively capture controlled, open-domain, and in-the-wild reasoning scenarios.
+consistently outperforms competitive baselines in both accuracy and efficiency.
+On the Vstar dataset, the -Patches variant achieves a 26.9% absolute accuracy gain, while on RealWorldQA, -Adaptive exceeds baseline performance by 12.1%.
+Moreover, in the TerraIncognita setting, our method reduces token usage by 67% while improving accuracy by 6.4%.
+These gains hold across multiple API providers, including GPT-4o111https://platform.openai.com/, Gemini-1.5Pro222https://gemini.google.com/, and Claude-3.5-Sonnet333https://anthropic.com/, demonstrating that generalizes well across model architectures and tokenization schemes.
+In summary, this work makes the following contributions: (1) We provide the first in-depth empirical analysis of how token allocation and spatial attention impact visual reasoning performance in black-box MLLMs.
+(2) We formulate the token-constrained, region-aware visual prompting problem and identify its core design dimensions.
+(3) We propose , a general-purpose framework for token-efficient, spatially coherent visual prompting.
+(4) We demonstrate that improves both accuracy and efficiency across models, datasets, and domains, highlighting a new path toward robust, resource-aware multimodal understanding.
 2 Pilot Experiments
-One of the primary challenges faced by black-box MLLMs is their inability to process visual inputs efficiently, leading to diminished accuracy in fine-grained visual tasks.
-Models like GPT-4o often struggle with recognizing detailed or occluded objects, particularly when dealing with complex images.
-These limitations are compounded by token constraints, which restrict the amount of image data that can be processed in a single prompt.
-Furthermore, according to the Vision pricing calculator444https://openai.com/api/pricing/, GPT handle images by resizing and splitting them into basic units of 512Γ512 pixels.
-Each of these units corresponds to 170 tokens.
-This method of processing not only imposes a strict limit on the image resolution but also increases the computational overhead due to the additional tokens generated from splitting.
-As a result, vital visual details may be lost when images are downsampled or resized to fit within the token limits, leading to poor performance on tasks that require precise visual grounding.
+A fundamental challenge in black-box multimodal large language models (MLLMs) is their inability to process high-resolution visual inputs efficiently under strict token budget constraints.
+In platforms such as GPT-4o, visual inputs are internally segmented into fixed-size image patches (typically 512Γ512 pixels), each corresponding to a predefined token cost (e.g., 170 tokens per patch)444https://openai.com/api/pricing/.
+This patch-based tokenization strategy, designed for balancing inference efficiency and fairness, introduces a critical trade-off: higher image resolution yields more detailed information but leads to rapid token consumption, limiting the number of patches a model can process.
+Conversely, reducing token usage through downsampling or cropping risks omitting essential fine-grained visual cues.
+To quantify the impact of this trade-off, we conducted a series of pilot experiments using GPT-4o-0513 on the Vstar-Bench dataset, which requires precise visual grounding of small or occluded objects in complex scenes.
+We compare three visual prompting strategies: (1) Unaltered Input, where the original image is submitted directly without preprocessing; (2) Image Crop, Unless otherwise stated, all regions of interest (ROIs) used in this paper are obtained automatically at inference time using the multi-scale, prompt-aware open-vocabulary detector pipeline described in SectionΒ LABEL:sec:emphasizer.
+We do not use any ground-truth bounding boxes from the datasets when constructing ROIs, including the introductory examples in this section; and (3) Zoomed Crop, which magnifies the cropped region to maximize visual feature visibility while preserving the 512Γ512 input constraint.
+Each cropped ROI is normalized to a consistent visual scale before composition.
+Concretely, we resize each ROI such that its longer side equals 512 pixels while preserving aspect ratio.
+ROIs that are originally smaller than this scale are not up-scaled beyond their native resolution, in order to avoid excessive blurring or hallucinations.
+For large input images, we first crop ROIs using the detector and then apply this max-sideβ512 rule per ROI.
+This normalization ensures a consistent scale before spatial composition in SectionΒ 4.2.
+The results in TableΒ 1 reveal several critical insights.
+While Unaltered Input provides full-scene context, it incurs excessive token cost (955 tokens), with limited performance gainsβsuggesting that high token usage alone does not guarantee better model understanding.
+In contrast, the Image Crop method reduces token usage substantially but fails to improve accuracy, likely due to the lack of visual emphasis and limited discriminative detail in the cropped region.
+The Zoomed Crop strategy significantly outperforms both baselines, achieving 64% accuracy with only 270 tokens.
+This improvement highlights the importance of visual emphasis: by magnifying relevant regions, Zoomed Crop restores critical visual granularity lost in standard downsampling, while maintaining input compactness.
+Importantly, this approach respects the modelβs fixed patch budget without requiring architectural changes or access to internal weights.
+These findings underscore a fundamental limitation in black-box MLLMs: existing visual prompting techniques fail to manage the resolutionβtoken trade-off effectively.
+NaΓ―ve approaches such as downscaling or region cropping without adaptive enhancement do not sufficiently preserve semantic or structural information.
+Our pilot experiments suggest that vision enhancement strategiesβparticularly those that incorporate selective magnificationβare essential for preserving task-relevant detail in constrained inference settings.
+In subsequent sections, we build upon this insight to design , a visual prompting framework that generalizes the principles of adaptive emphasis, spatial preservation, and budget-aware token allocation across diverse visual reasoning tasks.
 | Method | Accuracy |
@@
 | Zoomed Crop | 64% | 270 |
-To evaluate this issue, we conducted a series of pilot experiments using GPT-4o-0513 on the Vstar-Bench dataset.
-This dataset challenges MLLMs to accurately identify detailed objects within high-resolution images, making it an ideal test for the modelβs capacity to handle fine-grained visual information.
-The experiments compared three different image processing strategies: (1) an unprocessed prompt (Unaltered Input), where the image is fed to the model in its original form; (2) a prompt where the image is cropped to focus on the target object (Image Crop); and (3) a prompt where the cropped image is further enlarged to emphasize the most relevant visual features (Zoomed Crop).
-Both the Image Crop and Zoomed Crop methods were constrained to fit within GPT-4oβs patch size limit of 512x512 pixels.
-As shown in TableΒ 1, the Zoomed Crop method significantly outperformed the others, achieving an accuracy of 0.76 with a token usage of 270.
-In comparison, the Unaltered Input method, despite processing the entire image, only achieved an accuracy of 0.64 while consuming 955 tokens.
-Similarly, the Image Crop method, although reducing the token count to 270, did not yield any improvement in accuracy compared to the unprocessed input.
-These results highlight a fundamental problem in current black-box MLLMs: they fail to efficiently manage the trade-off between image resolution and token constraints.
-In cases like those presented by the Vstar-Bench dataset, where fine-grained visual information is critical, processing unaltered high-resolution images leads to excessive token consumption without improving accuracy.
-While the Image Crop method reduces token usage, it fails to improve performance because simply cropping an image without emphasizing key details does not provide sufficient context for the model to interpret the visual input accurately.
-The superior performance of the Zoomed Crop method underscores the importance of vision enhancement techniques in black-box MLLMs.
-By focusing on the most relevant portions of the image, Zoomed Crop preserves critical details while remaining within token limits, enabling the model to interpret detailed visual inputs more effectively.
-This approach resolves a common issue faced by black-box MLLMs, where downscaling or cropping images to meet token requirements often leads to a loss of essential information, reducing the overall effectiveness of the model.
-Our experiments reveal that without adaptive techniques like Zoomed Crop, black-box MLLMs struggle to process high-resolution images efficiently, limiting their performance on tasks that require precise visual recognition.
-These findings demonstrate the necessity of vision enhancement strategies to address the inherent limitations of token-constrained MLLMs.
 3 Related Work
@@
 These models rely on effective visual encoding strategies to bridge the gap between language and vision.
-Approaches such as CLIPΒ [31] align visual and language embeddings through contrastive learning, while models like FlamingoΒ [1] and BLIP-2Β [6] use cross-attention mechanisms or pretraining modules to link vision encoders with LLMs.
-However, these methods often rely on fixed low-resolution inputs (e.g., 224x224), limiting their ability to process high-resolution images or non-standard aspect ratiosΒ [16], which hampers performance on fine-grained tasks such as OCR and small object detection
-In contrast, open-source multimodal modelsΒ [15, 30, 36, 12, 37] allow for architectural modifications and fine-tuning to accommodate any-resolution inputs.
+Approaches such as CLIPΒ (yangAttentiveMaskCLIP2023) align visual and language embeddings through contrastive learning, while models like FlamingoΒ (alayracFlamingoVisualLanguage2022) and BLIP-2Β (daiInstructBLIPGeneralpurposeVisionLanguage2023) use cross-attention mechanisms or pretraining modules to link vision encoders with LLMs.
+However, these methods often rely on fixed low-resolution inputs (e.g., 224x224), limiting their ability to process high-resolution images or non-standard aspect ratiosΒ (liuVisualInstructionTuning2023), which hampers performance on fine-grained tasks such as OCR and small object detection
+In contrast, open-source multimodal modelsΒ (liLLaMAVIDImageWorth; xuLLaVAUHDLMMPerceiving2024; zhangLLaVAHDDivingHighResolution2024; liLLaVANeXTInterleaveTacklingMultiimage2024; zhaoMGLLaVAMultiGranularityVisual2024) allow for architectural modifications and fine-tuning to accommodate any-resolution inputs.
 However, black-box MLLMs such as GPT-4o and Gemini Pro, which impose strict token limits for computational efficiency, require alternative solutions.
 The need to downsample or crop images to meet these constraints often results in the loss of crucial visual details, particularly in tasks requiring detailed visual understanding.
-While position embedding interpolationΒ [2, 26, 19, 9, 5] and patch-based croppingΒ [30, 12] widely adpoted in open-soure models offer promising directions for any aspect ratio and any-resolution image processing, they are not applicable to black-box models, where architectural changes and extra training/fine-tuning are not permitted.
+While position embedding interpolationΒ (baiQwenVLVersatileVisionLanguage2023; wangQwen2VLEnhancingVisionLanguage2024; luoFeastYourEyes2024; hongCogAgentVisualLanguage2023; chenMiniGPTv2LargeLanguage2023) and patch-based croppingΒ (xuLLaVAUHDLMMPerceiving2024; liLLaVANeXTInterleaveTacklingMultiimage2024) widely adpoted in open-soure models offer promising directions for any aspect ratio and any-resolution image processing, they are not applicable to black-box models, where architectural changes and extra training/fine-tuning are not permitted.
 3.2 Object Detection
-Traditional object detection models, such as Faster R-CNNΒ [22] and YOLOΒ [21], effectively identify and localize objects within predefined categories.
+Traditional object detection models, such as Faster R-CNNΒ (renFasterRCNNRealTime2016) and YOLOΒ (redmonYouOnlyLook2016), effectively identify and localize objects within predefined categories.
 However, they struggle with open-set scenarios, where novel objects not seen during training need to be detected.
 Recent advances address this limitation through open-set detection models that leverage natural language processing.
-For instance, OV-DETRΒ [34] integrates CLIP with object detection to generate category-specific bounding boxes from textual prompts, enabling detection in open-world settings.
-Similarly, GLIPΒ [14] reframes detection as a grounding problem, improving alignment between visual regions and textual descriptions.
-DetCLIPΒ [32] extends this further using pseudo labels from large-scale captioning datasets, enhancing generalization.
-Grounding DINOΒ [17], built on the DETR frameworkΒ [3], also advances open-set detection through natural language integration.
-In addition, SAMΒ [11] and SAM-2Β [20] offer zero-prompt or minimal-prompt segmentation for arbitrary objects but lack robust text-prompt handling.
-EVF-SAMΒ [35] overcomes this by extending SAMβs capabilities to better manage complex text-based object segmentation.
+For instance, OV-DETRΒ (zangOpenVocabularyDETRConditional2022) integrates CLIP with object detection to generate category-specific bounding boxes from textual prompts, enabling detection in open-world settings.
+Similarly, GLIPΒ (liGroundedLanguageImagePretraining2022) reframes detection as a grounding problem, improving alignment between visual regions and textual descriptions.
+DetCLIPΒ (yaoDetCLIPDictionaryEnrichedVisualConcept2022) extends this further using pseudo labels from large-scale captioning datasets, enhancing generalization.
+Grounding DINOΒ (liuGroundingDINOMarrying2023), built on the DETR frameworkΒ (carionEndEndObjectDetection2020), also advances open-set detection through natural language integration.
+In addition, SAMΒ (kirillov2023segment) and SAM-2Β (ravi2024sam2) offer zero-prompt or minimal-prompt segmentation for arbitrary objects but lack robust text-prompt handling.
+EVF-SAMΒ (zhangEVFSAMEarlyVisionLanguage2021) overcomes this by extending SAMβs capabilities to better manage complex text-based object segmentation.
 By incorporating these models, enhances its ability to dynamically detect and emphasize regions of interest (RoIs), enabling black-box MLLMs to focus on the most relevant visual content without losing critical details, which is essential for maintaining high performance across varied resolutions.
 4 Method Overview
-Inspired by the observation derived from our pilot experiments, we propose , a comprehensive visual prompting mechanism designed to effectively address the loss of detail in images that occurs during the naive resizing process in current black box multimodal LLMs, such as GPT-4 and Gemini 1.5.
-As illustrated in FigureΒ 3, our mechanism comprises three key components: (1) A prompt-aware visual emphasizer that allocates high-fidelity image slices based on prompt texts to facilitate efficient and focused visual encoding; (2) A spatial-preserving encoding schema that consolidates the collected image slices while maintaining their relative spatial positions to create a condensed visual input; (3) A budget-aware prompting strategy that maximizes the accuracy of results obtained from the black box models while fits the budget requirement from users.
+Building upon the findings of our pilot experiments, we introduce , a unified visual prompting framework that enables detail-preserving, token-efficient encoding of high-resolution visual inputs for black-box multimodal LLMs.
+Existing models such as GPT-4o and Gemini 1.5 often rely on uniform downsampling or tiling to fit image inputs into fixed-size patches, which leads to significant information loss and degraded performance on fine-grained vision tasks.
+addresses this challenge through three tightly coupled modules, as shown in FigureΒ 3:
+β’
+Prompt-Aware Visual Emphasizer: extracts task-relevant image regions based on semantic cues derived from natural language prompts.
+β’
+Spatial-Preserving Orchestration Schema: reconstructs extracted regions into spatially faithful layouts that preserve object relationships and scene structure.
+β’
+Budget-Aware Prompting Strategy: manages token allocation under user-specified constraints by adapting the number and organization of image regions.
 4.1 Prompt-aware Visual Emphasizer
@@
 By eliminating these non-essential words, we concentrate on key terms that directly influence the visual emphasis.
-In addition to basic stopword removal, we utilize dependency parsingΒ [23, 7] to analyze the syntactic structure of the prompt.
+In addition to basic stopword removal, we utilize dependency parsingΒ (parse1; parse2) to analyze the syntactic structure of the prompt.
+We use a lightweight dependency parser to identify noun phrases and their syntactic roles (e.g., question objects versus modifiers).
+Rather than constructing a full scene graph, we use these dependencies as an importance-ranking mechanism: (1) higher-priority noun phrases are passed to the open-vocabulary detector first, and (2) when multiple entities compete for limited detection budget, the ranking is used to break ties.
+In other words, dependency parsing is used for importance weighting and ordering prior to detection, not for explicit relation modeling.
 This deeper analysis identifies core entities and relationships, such as subject-object pairs and action verbs, which are crucial for interpreting the userβs intent.
@@
 Multi-Scale Emphasizing Algorithm: Given a key object term extracted from the text prompt, the Multi-Scale Emphasizing AlgorithmΒ 1 utilizes a state-of-the-art object detection model to localize the corresponding object in the image prompt.
-In our experiments, we primarily employ GroundingDINOΒ [17] as our localization model.
+In our experiments, we primarily employ GroundingDINOΒ (liuGroundingDINOMarrying2023) as our localization model.
 The encoder in such models typically downsamples the input image to a resolution of or , potentially resulting in information loss when localizing the target object at a coarse granularity.
 To address this limitation, we propose a Multi-Scale Emphasizing Algorithm that processes the original image at multiple resolutions.
-The algorithm divides the input image into patches at various granularities, e.g.
-, , and beyond.
+The algorithm divides the input image into patches at various granularities, e.g., , , and beyond.
 For each generated patch, we apply the object detection model to localize the target object.
@@
 Image Shrinking: The resulting image, containing only the selected slices in their original positions with the rest left blank, is then shrunk to a predetermined size while maintaining its aspect ratio.
+In all configurations, Zoomer composes the selected ROIs and optional global view into a single image canvas, which is then passed to the black-box MLLM.
+In particular, in Zoomer-Adaptive and Zoomer-Global, ROIs are placed according to their original relative coordinates using the spatial-preserving layout, and, when enabled, a down-scaled global overview is appended as an additional strip on the same canvas.
+Thus, the MLLM always receives exactly one reconstructed image rather than multiple separate images.
 4.3 Budget-aware Prompting Strategy:
@@
 It provides the most detailed information but requires the largest token budget.
+Adaptive global view decision.
+Let EQΒ 1 denote the fraction of the original image covered by the selected ROIs.
+Zoomer-Adaptive decides whether to append a global overview based on : if , we append a down-scaled global view to preserve overall scene context; if , we omit the global view to save visual tokens, since the ROIs already cover a sufficiently large portion of the scene.
+The threshold is chosen empirically on the VSTAR validation set to balance accuracy and token efficiency.
 The selection among these varieties depends on the user-specified budget and the nature of the task.
@@
 | -Adaptive | 83% | 88% | 315 | 5112 | 3.1s | 9.8s | 1.575 | 25.56 |
+We evaluate the performance of on a diverse set of visual reasoning tasks using multiple black-box MLLMs.
+Our experiments are designed to assess both quantitative gains in accuracy and qualitative improvements in visual fidelity under constrained token budgets.
+Specifically, our evaluation is guided by the following research questions: RQ1: Accuracy.
+Does improve task accuracy across different black-box MLLMs on image-grounded reasoning tasks, compared to standard prompting baselines?
+RQ2: Token Efficiency.
+How does balance the trade-off between accuracy and token usage, and how does it perform relative to methods that either over-utilize or under-utilize the token budget?
+RQ3: Component Effectiveness.
+What is the contribution of each major component of βspecifically, the multi-scale visual emphasis module and the budget-aware prompting strategyβto the overall performance?
+Backbone MLLMs.
+We evaluate Zoomer on both commercial APIs (e.g., GPT-4o, Gemini, Claude) and open-source MLLMs.
+For the latter, we use Qwen2-VL-7B and InternVL2.5-8B as representative models in this work, primarily due to time and resource constraints.
+As shown in TableΒ 6, applying Zoomer on top of both open-source backbones consistently improves performance across four benchmarks, indicating that the proposed adaptive visual token allocation is not tied to proprietary APIs.
+Extending the evaluation to additional high-resolution open-source models such as Qwen3-VL, Qwen2.5-VL, LLaVA-UHD, and MG-LLaVA is an important direction for future work.
+5.1 Setup
+Assessment and Datasets We evaluated our system on a series of challenging multimodal tasks, using commercial black-box MLLMs for applications ranging from visual-language reasoning to image understanding and question answering.
+The experiments were conducted on a variety of different public datasets, including:
+1) VstarΒ (vstar): A benchmark dataset focused on image classification, used to evaluate fine-grained visual recognition capabilities in object detection and classification tasks.
+2) CVBenchΒ (tong2024cambrian1): Contains 2 sub-category, and , respectively, representing two-dimensional and three-dimensional visual image, respectively, to evaluate the performance of the model when processing images of different dimensions, especially the understanding ability in complex scenes.
+3) RealworldQAΒ (xAI2024): Used to test the multimodal question answering performance of the model in real-world scenarios, involving cross-language and cross-image information processing.
+4) MMVPΒ (mmvp): A validation set for multimodal visual processing, designed to evaluate the comprehensive understanding of models for complex visual scenes.
+5) ScienceQAΒ (sqa): A multimodal scientific question-answering dataset featuring multiple-choice questions across a diverse range of science topics.
+6) MMMUΒ yue2023mmmu: The validation part of a new benchmark, which designed to evaluate the performance of multimodal models on multidisciplinary tasks that require university-level subject knowledge and deliberate reasoning.
+7) HRΒ wang2024divideconquercombinetrainingfree: A high-resolution multimodal benchmark consisting of 4K and 8K images and corresponding questions.
 | API | Method | Vstar | CVBench-2D | RealworldQA | MMVP |
@@
 | Patches | 71.7% | 74.6% | 85.8% | 75.8% | 92.8% | 88.4% | 58.9% |
-In this section, we evaluate the performance of through a series of experiments designed to test its ability to improve token efficiency and preserve visual fidelity across different black-box MLLM.
-Specifically, we aim to answer the following questions: (i) Accuuracy: Does improve accuracy across different black-box MLLMs on image-related tasks?
-(ii) Efficiency: How does perform compared to baseline methods in terms of both accuracy and token efficiency?
-(iii) Component Contribution: What is the impact of key components in , such as multi-scale vision emphasize and the budget-aware prompt strategy?
-5.1 Setup
-Assessment and Datasets We evaluated our system on a series of challenging multimodal tasks, using commercial black-box MLLMs for applications ranging from visual-language reasoning to image understanding and question answering.
-The experiments were conducted on a variety of different public datasets, including:
-1) VstarΒ [28]: A benchmark dataset focused on image classification, used to evaluate fine-grained visual recognition capabilities in object detection and classification tasks.
-2) CVBenchΒ [24]: Contains 2 sub-category, and , respectively, representing two-dimensional and three-dimensional visual image, respectively, to evaluate the performance of the model when processing images of different dimensions, especially the understanding ability in complex scenes.
-3) RealworldQAΒ [29]: Used to test the multimodal question answering performance of the model in real-world scenarios, involving cross-language and cross-image information processing.
-4) MMVPΒ [25]: A validation set for multimodal visual processing, designed to evaluate the comprehensive understanding of models for complex visual scenes.
-5) ScienceQAΒ [18]: A multimodal scientific question-answering dataset featuring multiple-choice questions across a diverse range of science topics.
-6) MMMUΒ [33]: The validation part of a new benchmark, which designed to evaluate the performance of multimodal models on multidisciplinary tasks that require university-level subject knowledge and deliberate reasoning.
-7) HRΒ [27]: A high-resolution multimodal benchmark consisting of 4K and 8K images and corresponding questions.
 Models We employed three black-box MLLMsβGPT-4o-0513, Claude-v3-Sonnet, and Gemini-Proβaccessed via their respective APIs (OpenAI, Claude, Google).
@@
 Additionally, we compared token usage for each model configuration to evaluate the efficiency improvements offered by .
+Token efficiency metrics.
+Different providers adopt different image tokenization schemes.
+Whenever an API exposes image-token usage, we directly report the providerβs visual token count.
+When this information is not exposed, we instead count the number of patches that are actually transmitted after reconstruction as a proxy for visual tokens.
+To ensure fair cross-model comparison, we always measure token efficiency as a relative reduction with respect to each modelβs own Raw baseline. For a model and method , we compute
+and pair this with the corresponding accuracy change. This normalization prevents differences in provider-specific tokenization rules from biasing our conclusions about token savings.
 Baselines We compare against the following baseline methods:
@@
 5.2 Main results
-TableΒ 2 compares the performance of against baseline methods across various datasets using GPT-4o.
-The results show that , particularly in its Patches and Adaptive versions, consistently outperformed baseline approaches in terms of accuracy while maintaining lower token usage than the Raw method.
-consistently outperformed baseline methods, showing accuracy improvements up to 26% across multiple tasks.
-For example, in the Vstar dataset, -Patches achieved an accuracy of 0.717, compared to 0.565 using the Raw baseline, marking a 26.9% improvement.
-In RealworldQA, which demands complex multimodal reasoning, -Adaptive achieved 0.758 accuracy, outperforming the 0.676 accuracy of the Raw method by 12.1%.
-These results highlight that effectively preserves fine-grained visual details, enabling improved object recognition and image understanding across real-world tasks, where precise detail retention is crucial.
-We further evaluated on Claude-3.5-Sonnet and Gemini-1.5Pro to assess its generalizability across different black-box MLLMs.
-TableΒ 4 shows that demonstrated robust performance across different black-box MLLMs, for example, achieved an accuracy of 0.704 on Vstar, compared to 0.531 with the Raw baseline, marking a 32.6% improvement.
-Similarly, in Claude-3.5-Sonnet, outperformed the baseline by 34.5% on RealworldQA, improving from 0.610 to 0.741.
-These results suggest that can consistently enhance performance across different architectures, making it a versatile tool for various MLLM-based applications.
-A key contribution of is its ability to reduce token consumption while maintaining or improving accuracy.
-As shown in TableΒ 3, consistently delivers both token efficiency and performance improvements across various benchmark datasets.
-For instance, on the TerraIncognita dataset, guided by ManyICLΒ [10], achieves 0.83 accuracy using 315 tokens, compared to the Raw baselineβs 0.78 accuracy with 963 tokensβa 67% reduction in token usage while improving performance by 6.4%.
-Additionally, reduces latency, making it practical for real-time applications.
-In the TerraIncognita zero-shot setting, lowered latency from 4.8s to 3.1s, a 35.4% reduction without sacrificing accuracy.
-This makes highly suitable for tasks like autonomous driving, and real-time visual analytics, where both token efficiency and reduced latency are critical.
+RQ1: Accuracy Under Token Constraints
+We first evaluate the accuracy of across multiple visual reasoning benchmarks using GPT-4o.
+TableΒ 2 compares its performance with baseline prompting strategies, including Raw (unaltered image input), Crop, and heuristic patching.
+Across all datasets, demonstrates consistent and significant gains.
+Notably, -Patches achieves 71.7% accuracy on the Vstar dataset, outperforming the Raw baseline by 26.9% (from 0.565 to 0.717).
+On the more challenging RealWorldQA dataset, which requires complex visual-linguistic reasoning, -Adaptive yields 0.758 accuracyβ12.1% higher than the Raw input (0.676).
+These results indicate that enhances the modelβs ability to ground object references and interpret fine-grained visual cues by preserving high-fidelity regions while respecting token limits.
+RQ2: Cross-Model Generalizability
+To assess generalizability, we apply to other closed-source MLLMs, including Claude-3.5-Sonnet and Gemini-1.5Pro.
+As shown in TableΒ 4, the performance trends hold across architectures.
+For instance, on Vstar, improves accuracy from 0.531 to 0.704 on Gemini-1.5Proβa 32.6% gain.
+On RealWorldQA, Claude-3.5-Sonnet achieves a 34.5% improvement using (from 0.610 to 0.741).
+These results suggest that is model-agnostic and portable across different commercial MLLMs without requiring model access or retraining.
+As suggested by reviewers, we also evaluate Zoomer on open-source MLLMs.
+In particular, we use Qwen2-VL-7B and InternVL2.5-8B as representative backbones.
+As shown in TableΒ 6, applying Zoomer on top of both models yields consistent gains across all four benchmarks (e.g., +0.034β0.043 on RealWorldQA and +0.064β0.082 on VSTAR), confirming that our adaptive image focus mechanism is not tied to proprietary APIs.
+| Qwen2-VL-7B | InternVL2.5-8B | Qwen2-VL-7B + Zoomer | InternVL2.5-8B + Zoomer | |
+| CVBench-2D | 0.648 | 0.656 | 0.683 (+0.035) | 0.697 (+0.041) |
+| CVBench-3D | 0.632 | 0.629 | 0.675 (+0.043) | 0.669 (+0.040) |
+| RealWorldQA | 0.698 | 0.698 | 0.732 (+0.034) | 0.741 (+0.043) |
+| VSTAR | 0.584 | 0.573 | 0.648 (+0.064) | 0.655 (+0.082) |
+RQ3: Token Efficiency and Latency
+One of the central goals of is to maximize information utility under token constraints.
+TableΒ 3 reports token usage and performance on the TerraIncognita dataset under the ManyICLΒ (manyICL) setting.
+achieves 0.83 accuracy with 315 tokens, while the Raw method reaches only 0.78 with 963 tokensβyielding a 6.4% accuracy improvement with 67% fewer tokens.
+This substantial reduction in visual token cost translates directly into lower API usage and faster inference.
+Moreover, in real-time settings such as autonomous navigation, latency becomes critical.
+In a zero-shot TerraIncognita setting, reduces response time from 4.8 seconds to 3.1 seconds (a 35.4% reduction) without sacrificing accuracy.
+These results underscore the practical benefits of for latency-sensitive and resource-constrained applications such as real-time surveillance, embodied agents, and mobile vision systems.
+We further compare Zoomer against a stronger high-resolution baseline Vβ, which already uses high-resolution visual inputs without our prompt-aware cropping and composition.
+As shown in TableΒ 7, Zoomer still provides clear gains, especially on CVBench-3D and RealWorldQA.
+This indicates that adaptive image focus and spatial-preserving orchestration remain beneficial even when the baseline already operates in a high-resolution regime, and strengthens our connection to high-resolution prompting methods such as LLaVA-UHD and MG-LLaVA.
+| Method | VSTAR | CVBench-2D | CVBench-3D | RealWorldQA |
+| Zoomer | 0.717(+0.013) | 0.746(+0.045) | 0.883(+0.090) | 0.758(+0.060) |
+| Vβ | 0.704 | 0.701 | 0.793 | 0.698 |
 5.3 Findings
-Here we analyze why the -Patches version underperforms compared to -Global and even the Local version on certain datasets.
-For example, on the dataset, the accuracy of the Patches version is 0.025 lower than the Global version and 0.04 lower than the Local version.
-Similarly, on the MMVP dataset, the Patches version falls short by 0.005 compared to the Global version and by 0.003 compared to the Adaptive version.
-Given that these results are averaged over multiple measurements, and accounting for model fluctuations, we hypothesize that this performance drop occurs because the Patches version treats each RoI as an independent image and provides them separately to the MLLM.
-When there are too many RoIs, the model may fail to capture or integrate some of them, leading to a drop in accuracy.
+While -Patches generally performs well under generous token budgets, we observe that in certain datasets it underperforms relative to other variants.
+For instance, on the dataset, -Patches yields an accuracy that is 2.5 percentage points lower than -Global and 4.0 points lower than -Local.
+Similarly, on the MMVP benchmark, the Patches variant trails -Global by 0.5 percentage points and -Adaptive by 0.3 points.
+These differences persist across multiple trials and are consistent beyond minor fluctuations attributable to model non-determinism.
+We hypothesize that this degradation stems from the way -Patches encodes visual inputs.
+Specifically, each RoI is treated as an independent image patch and passed separately to the black-box MLLM.
+This approach, while maximizing local detail preservation, inherently disrupts the global spatial context and inhibits inter-region integration.
+When the number of RoIs increases, the model is forced to reason over fragmented visual inputs without spatial continuity, which may lead to failures in tasks requiring holistic scene understanding or relational reasoning.
+Failure case: missed small object in ROI extraction.
+Despite the overall gains, Zoomer may fail when the ROI extractor misses a small but question-critical object.
+FigureΒ 6 shows an example where the question asks: βIs the dog on the left or right side of the golden tower?β The dog is a tiny object in the scene.
+If the detector fails to propose an ROI covering the dog, Zoomer-Local will not include the dog in the reconstructed input, causing the MLLM to answer incorrectly.
+Zoomer-Adaptive may partially mitigate this issue by appending a global view, but it still cannot recover the missing evidence when the dog becomes indistinguishable in the down-scaled overview.
+This highlights a key limitation of detector-guided visual focus and motivates future detector-free visual focusing strategies.
+These findings suggest that preserving global structureβeven at the cost of slightly reduced local detailβis beneficial in tasks that involve complex spatial or semantic dependencies.
+They also underscore the need for future designs to better balance granularity with contextual coherence in visual prompting pipelines.
 5.4 Ablation Study
-To further understand the impact of the individual components within , we conducted an ablation study focusing on two key variants: 1) with multi-scale emphasize: Compared with the commonly used multi-resolution and directly use, multi-scale visual emphasis is used to identify and emphasize RoI in the image.
-2) with different models.
-To further investigate the contributions of βs components, we conducted an ablation study (TableΒ 5).
-The results demonstrate that the combination of multi-scale visual emphasis and Patches prompt strategies delivers the best performance across all most datasets.
-Comparing different vision emphasis models, such as EVF-SAM and Ground Dino, further highlights the effectiveness of .
-Despite differences in model capabilities, both show accuracy improvements across datasets.
-Additionally, when comparing different emphasis methodsβDefault, Multi-Resolution, and Multi-Scaleβthe Multi-Scale method consistently outperformed Multi-Resolution.
-We hypothesize that, while Multi-Scale crops images and may split objects, its pyramid-shaped multi-recall strategy compensates for this by enhancing recall.
-In contrast, although Multi-Resolution maintains object integrity, adjusting resolution disrupts the modelβs performance, likely because most models are trained on fixed-size inputs, and changing the resolution weakens their inherent capabilities.
+To assess the contribution of individual components in , we conducted a comprehensive ablation study along two primary dimensions: (1) the role of multi-scale visual emphasis strategies, and (2) the impact of different RoI localization models.
+The results are summarized in TableΒ 5.
+Patch selection under a fixed budget.
+As a stronger baseline than Raw and Resize, we compare Zoomer-Patches with Random and Uniform patch selection under the same patch budget (TableΒ 8).
+Since all three strategies transmit the same number of patches, the improvements of Zoomer-Patches can only come from semantically guided region selection and spatial-preserving composition, rather than from simply increasing visual resolution.
+| Method | CVBench-2D | CVBench-3D | VSTAR |
+| Zoomer-Patches | 0.746 | 0.858 | 0.717 |
+| Random-Patches | 0.571(-0.175) | 0.613(-0.245) | 0.597(-0.120) |
+| Uniform-Patches | 0.637(-0.109) | 0.731(-0.127) | 0.664(-0.047) |
+Visual Emphasis Strategies.
+We compare three emphasis methods: (i) a baseline strategy using the full image without emphasis ("Default"), (ii) a multi-resolution method that applies image resizing at various scales without cropping, and (iii) our proposed multi-scale strategy, which extracts overlapping RoI patches across hierarchical granularities.
+Across all tested datasets, the multi-scale approach consistently outperforms the other two.
+For instance, on the Vstar and CVBench datasets, the multi-scale strategy yields relative improvements of 8.2% and 6.4% over the multi-resolution method, respectively.
+This suggests that the hierarchical region recall provided by the multi-scale cropping mechanism compensates for potential object boundary fragmentation.
+In contrast, multi-resolution methodsβthough preserving global structureβnegatively impact performance, likely due to the modelβs reliance on fixed-size input distributions during pretraining.
+Adjusting resolution without architectural adaptation may lead to degraded feature extraction in the frozen backbone of black-box MLLMs.
+Backbone Variants for Region Extraction.
+We also evaluate the effect of using different visual backbone models for region proposal and saliency detection.
+Specifically, we compare GroundingDINOΒ (liuGroundingDINOMarrying2023) and EVF-SAM as RoI extractors in the multi-scale visual emphasis pipeline.
+Both variants yield significant improvements over the baseline; however, GroundingDINO exhibits slightly higher accuracy across datasets.
+This result suggests that while both models effectively localize task-relevant regions, the localization precision and language grounding of GroundingDINO may offer a marginal advantage in aligning visual slices with text prompts.
+Area Ratio Threshold
+We vary the threshold in Eq.Β equationΒ 1 among and report results on the VSTAR validation set.
+| Accuracy | Relative token usage | |
+| 0.25 | 0.645 | 0.82 |
+| 0.35 | 0.652 | 0.79 |
+| 0.45 | 0.647 | 0.76 |
+We observe that achieves the best trade-off between accuracy and token efficiency: smaller thresholds append the global view too often, increasing token usage with marginal accuracy gains, while larger thresholds skip the global view in cases that benefit from additional context.
+Synergistic Effects.
+Combining the multi-scale emphasis with the Patches variant of yields the strongest overall performance, validating the importance of both region-level granularity and architectural alignment.
+These findings highlight that visual prompting should be co-designed with content-aware pre-processing and prompt-space organization, especially under token-constrained inference regimes.
 6 Conclusion
-In this paper, we introduced , a novel visual prompting mechanism designed to overcome the limitations of black-box MLLMs in processing images while adhering to token constraints.
-Our approach effectively balances the need to capture essential visual details without exceeding token budgets, a challenge commonly encountered in existing models like GPT-4o and Gemini Pro.
-Through a comprehensive evaluation across datasets such as Vstar and RealWorldQA, demonstrated significant improvements, particularly in fine-grained visual tasks.
-Our results show that -Patches achieved a 26.9% accuracy gain over baseline methods in Vstar, and -Adaptive provided a 12.1% improvement in RealWorldQA.
-These gains were achieved while drastically reducing token usage, with delivering 6.4% higher accuracy in the TerraIncognita using 67% fewer tokens.
-Although this work primarily focuses on improving the efficiency of visual processing in black-box MLLMs, another potential issue that arises in real-world applications is communication cost.
-For example, transferring large images from edge devices (e.g., wearable cameras or glasses) to cloud servers can be expensive in terms of bandwidth, latency, and energy consumption.
-While is designed to reduce token usage, its application in minimizing data transmission costs is an area that could be explored in future work.
-As part of future research, we plan to investigate how could be adapted for edge ML applications, enabling local processing on devices such as wearable cameras.
-This would allow for more efficient handling of visual inputs at the edge, reducing the need for extensive data transfers to the cloud.
-We aim to measure latency, power consumption, and the overall impact on system performance to assess the feasibility of applying in these scenarios.
-In summary, offers a practical solution for enhancing visual processing in constrained MLLMs and opens up new directions for future exploration.
+This paper presents , a novel visual prompting framework that addresses the challenge of preserving fine-grained visual detail under token constraints in black-box multimodal language models (MLLMs).
+By decomposing the prompting task into prompt-aware emphasis, spatially structured orchestration, and budget-aware input construction, enables effective visual grounding without access to model internals or architectural modifications.
+Comprehensive evaluations across datasets such as Vstar, RealWorldQA, and TerraIncognita demonstrate that consistently improves task accuracy while reducing token consumption.
+For example, -Patches achieves a 26.9% gain over baseline accuracy on Vstar, while -Adaptive improves RealWorldQA accuracy by 12.1%.
+On TerraIncognita, achieves higher accuracy with 67% fewer tokens, highlighting its practical utility in resource-constrained settings.
+While the present work focuses on token-level efficiency, an important emerging challenge is the communication cost associated with transferring high-resolution images from edge devices to cloud-based MLLMs.
+This concern is particularly relevant in scenarios such as wearable computing or mobile robotics.
+A promising direction for future research is to extend toward edge-aware visual prompting, enabling lightweight pre-processing directly on-device to reduce upstream bandwidth and latency while maintaining task performance.
+In summary, contributes a modular and model-agnostic framework for enhancing visual processing in black-box MLLMs.
+Its integration of spatial structure, semantic relevance, and token budget constraints offers a principled foundation for both immediate deployment and future extensions in bandwidth-constrained or real-time vision-language applications.
+From heuristics to explicit optimization.
+The current Zoomer pipeline is heuristic in nature: it relies on prompt-aware detection, spatial-preserving composition, and simple token budgeting rules.
+A natural next step is to formalize adaptive visual token allocation as a constrained optimization problem, e.g., maximizing task accuracy or expected utility under a fixed visual-token budget and latency constraint, and to develop trainable surrogates or reinforcement learning policies that optimize this objective end-to-end.
+Beyond external detectors.
+Zoomer currently relies on open-vocabulary detectors (e.g., GroundingDINO, EVF-SAM) applied in a multi-scale fashion, which can become a bottleneck when detector performance degrades.
+Our experiments partially mitigate this via Zoomer-Adaptive and Zoomer-Global, which automatically include a global view when the cropped area ratio is small.
+Nonetheless, exploring detector-light or detector-free mechanismsβsuch as leveraging internal attention maps or saliency predictions from open-source MLLMsβis an exciting direction for future work.